Grade 12th passGeneral Physics

What should be the minimum average velocity of water in a small tube of radius 5 mm, so that the flow is just turbulent? (Given: Viscosity of water=0.01 poise and density of water = 1000 kg-m⁻3)

UNNATI AGARWAL
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Here r = 0.005 m , diameter D =2r = 0.010 m
η = 0.001 Pa?s , ρ = 1000 kgm?³
For flow to be just turbulent, Re = 3000
∴  v = Re η/ρD = 3000 × 0.001/1000 × 0.010 = 0.3 ms?¹
∴ Viscous force on the drop, ( According to Stoke's Law)
F = 6πηrv
= 6 × 3.14 × 1.8 × 10?5 × 2 × 10?5 × 5.8 × 10?²
= 3.94 × 10?¹° N